OpenStudy (anonymous):

ALGEBRA 1! Need help immediately! Will fan and give medal to first answer. Question is MULTIPLE CHOICE!

OpenStudy (anonymous):

Solve the following system of equations: -2x + y = 1 -4x + y = -1 1. (3, 1) 2. (-1, 3) 3. (-1, -3) 4. (1, 3)

OpenStudy (anonymous):

You need to use substitution

OpenStudy (e.mccormick):

Or elimination. It would work well here with a sign flip on the second equation.

OpenStudy (anonymous):

Okay thanks guys!

OpenStudy (e.mccormick):

np. have fun!